9x+4y=-17
12y=-3-3x

Answer by checkley77(12844)

9x+4y=-17
12y=-3-3x or 3x+12y=-3 now multiply this equation by (-3) & add.
-9x-36y=9
9x+4y=-17
-----------------
-32y=-8
y=-8/-32
y=1/4 answer
9x+4*1/4=-17
9x+1=-17
9x=-17-1
9x=-18
x=-18/9
x=-2 answer.